OpenStudy (anonymous):

\[\lim_{z \rightarrow 0} \frac{1}{\sqrt[z^{2}]{e}}\]

OpenStudy (lgbasallote):

\[\Large \lim_{x \rightarrow 0} (\frac{1}{e^{1/x^2}})\]

OpenStudy (lgbasallote):

use l'hospital

OpenStudy (anonymous):

I learned limits for the first time this week...I don't know what l'hospital is

OpenStudy (lgbasallote):

hmm okay...uhmm...

OpenStudy (anonymous):

I don't understand how you can take a zeroth root of something

OpenStudy (anonymous):

Or a fractional root of something

OpenStudy (anonymous):

Hurts my head

OpenStudy (lgbasallote):

try substituting 0 into \[\frac{1}{e^{1/x^2}}\] what do you get?

OpenStudy (anonymous):

Undefined

OpenStudy (lgbasallote):

have you already discussed that \[\frac{1}{0} = \infty?\]

OpenStudy (anonymous):

No, I only know \frac{1}{0} is undefined

OpenStudy (anonymous):

Okay, this is making me believe I wrote down the wrong homework numbers and I'm doing the incorrect exercises...

OpenStudy (anonymous):

How can I solve this with methods we haven't learned in class yet

OpenStudy (lgbasallote):

well i dont know what you have learned do i

OpenStudy (mathteacher1729):

@lgbasallote \(1/0 \) is NOT defined. The LIMIT as x goes to zero of 1 / x is + infinity from the left (from numbers greater than zero) and - infinity from the right (numbers less than zero).

OpenStudy (mathteacher1729):

This is a very very very important notion in calculus. Evaluating a function at a point and taking the limit as it approaches that point are not always the same. :)

OpenStudy (lgbasallote):

hmm good point

OpenStudy (lgbasallote):

that's also the reason why i hate limits lol

OpenStudy (anonymous):

I thought limit values are not allowed to be infinity?

OpenStudy (mathteacher1729):

I highly recommend (to anyone studying limits) to read this excellent set of notes from PSU (my undergrad university) :) "relative rates of growth" PDF http://www.math.psu.edu/files/141rates1.pdf

OpenStudy (anonymous):

Thank you mathteacher1729